OpenStudy (anonymous):

Solving Special Systems Classify system. Give the number of solutions. y=2x-3 y-2x=-3

OpenStudy (anonymous):

@texaschic101 this ?

OpenStudy (texaschic101):

y = 2x - 3 y - 2x = -3 y = 2x - 3 same line...infinite solutions

OpenStudy (anonymous):

@texaschic101 3x+y=4 -3x=y-7